schedules downloaded and automatic loading of A1513 appended (Lucia)

Ops2 quite slow


module HOB+1011; starting position: 14584.546 GB

  • disk record failure when recording first scan. I swapped banks to other module. First recorded scan at 17:40. (scan 055-1739). DISK change will be necessary (Lucia)

Changed module to HOB+0100 (16TB) (Imogen)

Had to stop and restart schedule several times to ensure mk5hb was working correctly. No scans lost however, as recording on flexbuf did not stop. (Imogen)


module HOB+0124 (different from spreadsheet); starting position: 16085.055 GB


module HOB+0102 (different from spreadsheet); starting position: 1104.741 GB

  • DBBC restarted after R1 experiment has finished (power-cycle). Tsys look better now (Lucia)


  • cannot connect to ERC server. Wrote email to John. (Lucia)
  • 18:43 UT - Jon restarted the ERC server and we are monitoring now. (Lucia)